Can someone help me on this?
Is -4 a solution to the equation: 7x -5 + 3x = 6 + x - 10
Thanks,
Asley

Answer by stanbon(75887)

Is -4 a solution to the equation: 7x -5 + 3x = 6 + x - 10
If -4 is a solution then:
7(-4)-5+3(-4)=6+(-4)-10
-28-5-12=-8
-45=-8
It doesn't so -4 is not a solution.
Cheers,
Stan H.